OpenStudy (anonymous):

Find the balance in the account after the given period. $3500 deposit earning 6.75% compounded monthly, after 6 months.

OpenStudy (wolf1728):

First, you need a formula

OpenStudy (wolf1728):

Total = Principal * (1 + (.0675/12))^12*.5 Total = 3,500 * (1 + (.0675/12))^6 Total = 3,500 * (1 + (0.005625))^6 Total = 3,500 * (1.005625)^6 Total = 3,619.7986439863 Total = 3,619.80 (rounded)